A coint rests on a record 0.13 m from its center. The record turns on a turntable that rotates at variable speed. The coefficient of static friction between the coin and the record is 0.29. What is the maximum coin speed at which it does not slip?
Answer #1

Let the max speed be vmax.

Then centripetal force acting on coin, Fc=mv2/d

where, d=0.13m

Now, Force of fiction, Ff=?mg

For max speed without slipping:

Fc=Ff

mvmax2/d=? mg

vmax2/0.13=0.29*9.8

v=0.608 m/s